OpenStudy (anonymous):

Hi I need help with ALgebra 2 problem! Can you just give me a sec to write down the equation that would be great!

OpenStudy (anonymous):

\[3(x+3)^{3/4}= 81\]

OpenStudy (anonymous):

I can try I'm in 7th grade. I'm in Algebra as well. :)

OpenStudy (anonymous):

So these are my steps I will show okay give me one sec

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

First I divided the 3 from both sides to get rid of that number. I got \[(x+3)^{3/4}=27\]

OpenStudy (anonymous):

so then what I did was I did the opposite of the fraction sign.. and got \[\left| x+3 \right|=81\]

OpenStudy (anonymous):

and then what I did was got this as my answer \[x+3= plus or minus 81\]

OpenStudy (anonymous):

and then plugged 81 and -81 into x and got -78 and 84

OpenStudy (anonymous):

but then i looked online for the answer that my techer put up and he had it without the absolute value sign but then that confused me with the rule you use

OpenStudy (haleyelizabeth2017):

I have never done it that way before.....weird

OpenStudy (anonymous):

Hm, Yeah.

OpenStudy (anonymous):

yeah like it goes with this rule...but apparently I did it wrong but i followed the rules...

OpenStudy (haleyelizabeth2017):

You needed to subtract 3 from both sides before you plug it in...

OpenStudy (haleyelizabeth2017):

That is the only thing I can tell....

OpenStudy (anonymous):

oh yeah... dude thanks!

OpenStudy (haleyelizabeth2017):

Welcome ;)

OpenStudy (anonymous):

I agree... If you don't mind may I ask for help with a question as well? Which of the following has a graph that is a straight line? Equation 1: y = 4x3 + 6 Equation 2: y = 5x − 4.5 Equation 3: y2 = x − 1 Equation 4: y = 2x2 + 6 I'm confused. :(

OpenStudy (haleyelizabeth2017):

You might get more help if you post it in a new question ;)

OpenStudy (anonymous):

.... girl it came out all funky I cant see a single thing.

OpenStudy (anonymous):

yeah you should post a new one! Thanks guys!

OpenStudy (anonymous):

Okay, Thank you anyways! ;)

OpenStudy (anonymous):

your welcome!

OpenStudy (anonymous):

thanks both of you!

OpenStudy (haleyelizabeth2017):

:)